An emf source of 6.0V is connected to a purely resietive lamp and a current of 2.0 amperes flows. All the wires are resistance free. What is the resistance of the lamp?
1
Expert's answer
2018-03-31T09:38:08-0400
R = U/I = 6V/2A = 3 Ohm